An 8.5% coupon, 25 year, $1000 face value bond presently has a yield to maturity of 9.75%. Assuming annual interest payments, what is the price of the bond?
Answer #1

Annual coupon=1000*8.5%=85

Hence price of bond=Annual coupon*Present value of annuity factor(9.75%,25)+$1000*Present value of discounting factor(9.75%,25)

=85*9.25437694+$1000*0.097698248

=$884.32(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=85[1-(1.0975)^-25]/0.0975

=85*9.25437694

2.Present value of discounting factor=1000/1.0975^25

=1000*0.097698248